OpenStudy (anonymous):

Find the next three terms in the pattern. 10, 15, 25, 40, ….

OpenStudy (anonymous):

60, 85, 115

OpenStudy (welshfella):

@BweadedChicken - please explain how you got that answer

OpenStudy (anonymous):

^

OpenStudy (welshfella):

@fieldhockey00 - do you see the logic behind this ?

OpenStudy (anonymous):

it started by +5, but it increases the value of how much he added by +5 so its changing the value of how much hes adding, +5, +10, +15

OpenStudy (♪chibiterasu):

Notice the increment: 15 - 10 = 5 25 - 15 = 10 40 - 25 = 15

OpenStudy (♪chibiterasu):

The value in which each of the numbers are increased by is increasing by 5 itself after each number.

OpenStudy (anonymous):

Thank you! You both are life savers!!!! c: